MILLIMETER AND SUB-MILLIMETER SPECTROSCOPY OF CRCCH (X^{6}Σ^{+})

JIE MIN, LUCY. M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of CrCCH has been measured using millimeter/sub-millimeter direct absorption methods in the frequency range of 445-485 GHz. The molecule was created in an AC discharge of Cr(CO)_{6}, acetylene and argon. Six rotational transitions were measured. The spectra clearly indicate that the electronic ground state is a sextet. Spectra of CrCCD have also been recorded using the mixture of Cr(CO)_{6}, deuterated acetylene and argon in an AC discharge. The data were fit with a Hunds case (b) Hamiltonian and rotational, spin-rotation and spin-spin parameters have been determined. The structure has been calculated based on the rotational constants.